Solve each system using substition. Write no solution or infinitely many solutions where appropriate.
+y=+3x+-10+
+AND+
y=+2x+-5+
PLEASE AND THANK YOU!!

Answer by stanbon(75887)

Solve each system using substition. Write no solution or infinitely many solutions where appropriate.
+y=+3x+-10+
+AND+
y=+2x+-5+
-----------------------------
Since both expressions are equal to "y",
3x-10 = 2x-5
Subtract 2x from both sides to get:
x - 10 = -5
Add 10 to both sides:
x = 5
---
Since y = 2x-5, y = 2*5-5 = 10-5 = 5
-----
Solution:
x = 5 ; y = 5
=====================
Cheers,
Stan H.